An RLC series circuit has a 5.00Ω, a 20.00mH inductor, and a 30.00µF capacitor, and has a 110V source. a. Find the circuit’s impedance at 60 Hz. b. What is the current at the 60Hz frequency? c. What is the resonant frequency of the circuit? d. What is the average power dissipated at the circuit’s resonance frequency?
Answer #1

Solution : 3. for a RLC Series củwwt :- R=5.00st L=20.00 mH = 20.00XP H (= 30.00MF - 30.00x10-6F Vo = llov (a) f 2r ImpedanceInductive reactance (X) = wl Xi = 376.8 * 20.00 [X = 7.53652 x/03 52 Tmpedanu (z) ✓ R² + (x - Xe) 2 | (5.00) ² + + (7.536-88.Resonant Frequency (fo) = 2X3.14 / 20.00x10–3 x 30.00x/0-6 fo - 6:28 x 1600x16T X 60 -4 6.28 x 560X60-4 205.58H2 Ifo (d) At r